Application of Functions
1. If a sample of gas such as helium is​ heated, it will expand. The formula​
V(T) = 0.147T​ + 40 calculates the volume V in cubic inches of a sample of gas at temperature T in degrees Celsius.
Answer the following questions.
​a) Evaluate​ V(0) and interpret the results.
​V(0)=________
Interpret the​ results:
A. At 40 degrees Celsius the volume of the gas is 0 cubic inches.
B. At 0 degrees​ Celsius, the volume of the gas is 40 cubic inches.
b) What is the volume of gas when the temperature is 100 degrees​ Celsius?
 The volume is _______cubic inches.
 (Round your answer to the nearest​ tenth.)
​c) Write a formula that gives the volume of a gas that at 0 degrees Celsius has a Volume of 137 cubic centimeters. This gas increases in volume by 0.5 cubic centimeters for every 1 degree increase in temperature.
Fill in the values for the​ formula: V(T)​ = ___T+___

Answer by Boreal(15235)

v(0)=40 cu in, since the t term is 0.
0=0.147(0)+40 , which is 40.
40=0.147(40)+40, which is inconsistent. 40 does not equal 98.8; besides, there can be no volume of 0 for a gas at a temperature > absolute 0.
v(100)=0.147*100+40=14.7+40=54.7 cu in
the last one would be v(t)=0.5t+137, the rate of increase being the slope and the constant being the y-intercept.